### v2

# Qudit-native measurement protocol for dynamical correlations using Hadamard tests

Abstract
Dynamical correlations reveal important out-of-equilibrium properties of the underlying quantum many-body system, yet they are notoriously difficult to measure in experiments. While measurement protocols for dynamical correlations based on Hadamard tests for qubit quantum devices exist, they do not straightforwardly extend to qudits. Here, we propose a modified protocol to overcome this limitation by decomposing qudit observables into unitary operations that can be implemented and probed in a quantum circuit. We benchmark our algorithm numerically at the example of quench dynamics in a spin-1 XXZ chain with finite shot noise and demonstrate advantages in terms of signal-to-noise ratio over established protocols based on linear response. Our scheme can readily be implemented on various platforms and offers a wide range of applications like variational quantum optimization and probing thermalization in many-body systems.
I Introduction
The basic unit in which information is encoded in quantum computers is typically the qubit, the quantum generalisation of the classical bit. However, many physical systems naturally offer units with a higher-dimensional local Hilbert space. Embracing these higher dimensions leads to new ways for processing information based on qudits [1], with applications like alternative quantum algorithms [2], optimal measurements [3], and native encoding of complex quantum systems with higher local degrees of freedom [4]. On a fundamental level, the different coherence [5], dissipation, and entanglement structure [6] of qudit systems with respect to qubits can offer advantages for noise resilience [7] and quantum error correction [8]. Thanks to these favorable properties, qudits have gained traction on various quantum technology platforms [9, 10, 11] with manifold applications, e.g., in quantum cryptography [12, 13] and quantum simulation [14, 15, 16].
[…]
In this work, we formulate the Hadamard-test procedure for qudit systems on deterministic quantum circuits, enabling measurements of dynamical correlation functions in many-body systems with a higher-dimensional local Hilbert space. This is achieved by decomposing dynamical correlation functions of arbitrary observables into sums of unitaries, which can be measured with Hadamard tests on the quantum device, requiring only a single ancillary qubit. We numerically test the performance of our protocol by simulating a measurement of the two-time commutator and anti-commutator in a spin-1 XXZ chain with finite shot noise, where our scheme proves to be more efficient than established linear-response-based techniques [18, 19]. Our measurement protocol can readily be implemented on various qudit quantum devices, e.g., based on trapped ions, superconducting circuits, or neutral atoms, and thus provides a versatile tool for characterizing many-body physics on these platforms.
Figure 1: Quantum circuit for measuring dynamical correlations based on Hadamard tests, extended to qudit devices. The general structure consists of a qudit register that implements the physical system of interest plus an ancillary qubit. The register is initialized in state $|\psi_{0}\rangle$ and evolves under the unitary time-evolution operator $\hat{U}(t)$. The ancilla is subject to bit flip gates $\hat{X}$ and the Hadamard gate $\hat{H}$. In order to obtain the real and the imaginary parts of the dynamical correlation $\langle\hat{A}(t_{1})\hat{B}(t_{2})\rangle$, the ancillary qubit is initialized in the state $|\phi^{\pm}_{0}\rangle=(|0\rangle+e^{i\alpha_{\pm}}|1\rangle)/\sqrt{2}$ with $\alpha_{+}=0$ and $\alpha_{-}=\pi/2$, respectively. The decomposition of the qudit observables $\hat{A}$ and $\hat{B}$ into a sum of two unitaries $\hat{W}_{X}$ and $\hat{W}_{X}^{\dagger}$ with $X\in\{A,B\}$, makes it possible to implement their action as controlled quantum gates between the ancilla and the register. After the circuit is executed, measuring the probability of finding the ancilla in the state $|0\rangle$ gives access to the desired value of the dynamical correlation function.
This paper is organized as follows. In Section II, we present the general measurement protocol in a platform-agnostic framework. Section III puts forward our numerical results illustrating the Hadamard-test protocol for measuring dynamical correlations in a spin-1 XXZ chain with finite shot noise, followed by a comparison to linear-response-based schemes and a discussion of perspectives for an experimental realization. Finally, we present our conclusions in Section IV.
II Measurement protocol
Our goal is to measure dynamical correlation functions of two Hermitian (multi-)qudit observables $\hat{A}$ and $\hat{B}$ of the form
$C_{AB}(t_{1},t_{2})=\big{\langle}\hat{A}(t_{1})\hat{B}(t_{2})\big{\rangle}\,,$ (1)
where we use the shorthand notation $\hat{O}(t)=\hat{U}^{\dagger}(t)\hat{O}\hat{U}(t)$. The unitary operator $\hat{U}(t)$ may equally well describe a Hamiltonian time evolution (e.g., in cold-atom experiments) or a sequence of quantum gates on a digital quantum device, for example, representing a parametrized quantum circuit in the context of variational quantum simulation algorithms [38]. In general, the two-time correlator in Eq. 1 is a complex-valued quantity, $C_{AB}=C_{AB}^{+}/2-iC_{AB}^{-}/2$, whose real and imaginary parts are given, respectively, by the two-time anti-commutator
$C_{AB}^{+}(t_{1},t_{2})=\big{\langle}\big{\{}\hat{A}(t_{1}),\hat{B}(t_{2})\big{\}}\big{\rangle}\,$ (2a)
and commutator
$C_{AB}^{-}(t_{1},t_{2})=i\big{\langle}\big{[}\hat{A}(t_{1}),\hat{B}(t_{2})\big{]}\big{\rangle}\,.$ (2b)
Figure 1 illustrates the quantum circuit implementing the measurement protocol for dynamical correlations. The qudit register representing the physical system of interest is initialized in the state $|\psi_{0}\rangle$. Its evolution under the operator $\hat{U}(t)$ is interrupted at the times $t_{1}$ and $t_{2}$ by controlled gates coupling the register to an ancillary qubit. These gates represent the action of observables $\hat{A}$ and $\hat{B}$, whose dynamical correlation we intend to measure. The value of the corresponding correlation function is obtained by measuring the probability of the ancillary qubit to be in the computational $|0\rangle$ state at the end of the circuit. So far, the protocol corresponds to the standard Hadamard-test procedure well-known for qubits [38], which strongly relies on the fact that observables are both Hermitian and unitary. By contrast, for qudits, operators $\hat{A}$ and $\hat{B}$ are typically not unitary and therefore do not directly correspond to physical quantum gates.
In order to extend such a protocol to qudits, we represent $\hat{A}$ and $\hat{B}$ as a sum of unitary operations that can be inserted in the quantum circuit. The procedure can be seen as a special case of the more general linear combination of unitaries (LCU) method [54]. In general, we can decompose a Hermitian qudit operator $\hat{X}$ into a sum of a unitary operator $\hat{W}_{X}$ and its Hermitian conjugate as
$\hat{X}=\frac{1}{2}\big{\lVert}\hat{X}\big{\rVert}(\hat{W}_{X}+\hat{W}_{X}^{\dagger})\,,$ (3)
[…]
$\hat{W}_{X}=\dfrac{\hat{X}}{\lVert\hat{X}\rVert}+i\sqrt{\mathds{1}-\dfrac{\hat{X}^{2}}{\lVert\hat{X}\rVert^{2}}}\,.$ (4)
This decomposition allows us to rewrite Eq. 1 as a sum of correlation functions of unitaries,
$C_{AB}(t_{1},t_{2})=\frac{1}{4}\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}C_{V_{A}V_{B}^{\prime}}(t_{1},t_{2})\,.$ (5)
The real and imaginary parts of the correlators appearing on the right-hand side of Eq. 5 can now be accessed by distinct realizations of the circuit in Fig. 1, where the controlled gates correspond to the unitaries $\hat{W}_{A}^{(\dagger)}$ and $\hat{W}_{B}^{(\dagger)}$. Let $|\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\rangle$ be the total quantum state after execution of the circuit, given the initial ancilla state $|\phi_{0}^{\pm}\rangle=(|0\rangle+e^{i\alpha_{\pm}}|1\rangle)/\sqrt{2}$ with $\alpha_{+}=0$ and $\alpha_{-}=\pi/2$ and intermittent controlled unitaries $\hat{V}_{A}\in\{\hat{W}_{A},\hat{W}^{\dagger}_{A}\}$ and $\hat{V}_{B}\in\{\hat{W}_{B},\hat{W}^{\dagger}_{B}\}$. As shown in Appendix A, the value of the two-time (anti-)commutator can be extracted from the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})=|\langle 0|\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\rangle|^{2}$ of measuring the ancillary qubit in the state $|0\rangle$ according to
$C_{AB}^{\pm}(t_{1},t_{2})=\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\Big{(}P_{V_{A}V_{B}^{\prime}}^{\pm}(t_{1},t_{2})-\frac{1}{2}\Big{)}\,.$ (6)
Therefore, the same circuit allows one to measure either the real or the imaginary part of the dynamical correlation function simply by preparing the ancillary qubit in a different initial state.
A quantity of practical importance is the variance of the estimator for the dynamical correlation function in Eq. 5, as it determines the number of measurements that need to be performed on the quantum device to obtain $C^{\pm}_{AB}$ with sufficient precision. In Appendix B, we derive the result If the probabilities $P_{V_{A}V_{B}}^{\pm}$ are estimated by repeating the Hadamard test with controlled unitaries $\hat{V}_{A}$ and $\hat{V}_{B}$, respectively, $M_{V_{A}V_{B}}$ times, the two-time (anti-)commutator can be estimated with variance
$\operatorname{Var}[C^{\pm}_{AB}]=\big{\lVert}\hat{A}\big{\rVert}^{2}\big{\lVert}\hat{B}\big{\rVert}^{2}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\frac{P^{\pm}_{V_{A}V_{B}^{\prime}}\big{(}1-P^{\pm}_{V_{A}V_{B}^{\prime}}\big{)}}{M_{V_{A}V_{B}^{\prime}}}\,.$ (7)
The derivation of this result is reported Appendix B. Although the precise value of the variance ultimately depends on the details of the system, a useful a priori upper bound is given by the product of the square norms of observables $\hat{A}$ and $\hat{B}$, $\operatorname{Var}[C^{\pm}_{AB}]\leq M^{-1}\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$, assuming $M$ shots are allocated for measuring each configuration. This result is consistent with the bound on the variance directly derivable from Eq. 2 for a (hypothetical) projective measurement of the (anti-)commutator with a total budget of $4M$ shots, $\operatorname{Var}[C^{\pm}_{AB}]\leq(4M)^{-1}4\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$.
In what follows, we illustrate our protocol for measuring dynamical correlations using numerical simulations and benchmark its performance with finite measurement statistics.
III Numerical results
Figure 2: Dynamical correlation functions after a quench in the spin-1 XXZ chain model. The system of $N=10$ spins is initialized in the equal superposition of the two degenerate ground states of the model for $J_{z}/J_{xy}\to\infty$ and, after the quench, undergoes unitary evolution under the Hamiltonian $H_{0}$ with $J_{z}/J_{xy}=0.5$. The solid lines show the time trace of the two-time (a) anti-commutator and (b) commutator for the Hadamard-test protocol (blue) and for the linear-response protocol (red) with perturbation strength $\lambda=0.2$ and a pulse area $J_{xy}\Delta t=10^{-3}$. The shaded areas mark the standard error of the mean computed directly from the quantum state, while the colored circles correspond to the expectation values obtained from sampling with finite statistics, illustrating the effect of shot noise. The Hadamard-test measurement of the anti-commutator (commutator) was sampled with 1500 (8000) shots per point, whereas the sampled linear-response measurement corresponds to 1500 (12000) shots per point. For a comparable total budget of shots, the Hadamard-test protocol achieves a better signal-to-noise ratio than the linear-response protocol throughout the simulation.
In this section, we illustrate our protocol via numerical simulations for the example of quench dynamics in a spin-1 XXZ chain. This model represents a natural extension to qudit systems from its spin-$1/2$ version, which has served as a paradigmatic platform for testing novel quantum algorithms on qubit-based quantum hardware in the context of many-body quantum simulations [55]. Notably, closely related spin-1 models have been implemented on recently developed qudit quantum processors [56], representing promising targets for our proposed protocol. Furthermore, the spin-1 XXZ chain exhibits a richer phase diagram than its spin-$1/2$ counterpart, including the emergence of topological phases [57, 58, 59]. These properties make the model appealing for demonstrating and validating our qudit-native approach for measuring dynamical correlations.
Apart from statistical errors, our protocol is exact in the sense that it does not involve approximations that could lead to systematic errors. This is in contrast to protocols based on linear response, where by design a trade-off between accuracy and signal-to-noise ratio has to be made. We therefore benchmark our technique for measuring dynamical correlations based on Hadamard tests with a particular focus on its performance in a finite-statistics setting compared to linear-response measurements.
III.1 Quench dynamics in a spin-1 XXZ chain
We consider an XXZ chain of $N$ spin-1 particles, whose Hamiltonian is given by
$\hat{H}=\sum_{i=1}^{N-1}\left[J_{xy}\left(\hat{S}^{x}_{i}\hat{S}^{x}_{i+1}+\hat{S}^{y}_{i}\hat{S}^{y}_{i+1}\right)+J_{z}\hat{S}^{z}_{i}\hat{S}^{z}_{i+1}\right]\,.$ (8)
Here, $\hat{S}_{i}^{\alpha}$, with $\alpha\in\{x,y,z\}$, are spin-1 matrices acting on spin $i$, while $J_{xy}$ and $J_{z}$ denote the interaction strengths in the $xy$-plane and in $z$-direction, respectively. In order to initiate quench dynamics, we prepare the system deeply in the antiferromagnetic Néel phase, corresponding to the ground state for $J_{z}/J_{xy}\to\infty$. Since the ground state is degenerate in this limit, we choose a $\mathbb{Z}_{2}$-symmetric superposition of the two Néel states, but this choice is not crucial for our purposes. The system is then evolved under the Hamiltonian 8 with $J_{z}/J_{xy}=0.5$, which on the phase diagram lies in the bulk of the so-called Haldane phase [59]. While the initial state is characterized by antiferromagnetic order, which exhibits non-decaying spin–spin correlations, the Haldane phase is a gapped topological phase, characterized by non-vanishing string correlations. Moreover, it is considered a disordered phase; i.e., spin–spin correlations decay exponentially. Therefore, quenching from the antiferromagnetic phase to the Haldane phase, we expect to observe non-trivial dynamics in the two-point dynamical spin–spin correlations.
Specifically, we are interested in extracting the dynamical correlation functions in Eq. 2 for the $\hat{S}^{z}$ operator of two distinct spins $i$ and $j$ in the chain, denoted by $C^{\pm}_{ij}(t_{1},t_{2})\equiv C_{S_{i}^{z}S_{j}^{z}}^{\pm}(t_{1},t_{2})$. Following standard conventions [60], we consider only the “connected part” of the two-time anti-commutator, $C^{+}_{ij}(t_{1},t_{2})\to C^{+}_{ij}(t_{1},t_{2})-2\langle\hat{S}_{i}^{z}(t_{1})\rangle\langle\hat{S}_{j}^{z}(t_{2})\rangle$ (the product of expectation values forming the “disconnected part” can be obtained from standard projective measurements). In Fig. 2, we show the time trace of the two-time (anti-)commutator $C_{12}^{\pm}(0,t)$ after the quench.
III.2 Hadamard-test protocol with finite measurement statistics
In this section, we investigate how well our Hadamard-test protocol reproduces the above ideal results with finite statistics.
III.2.1 Implementation of the controlled unitaries for spin observables
In order to implement our protocol, we need to calculate the explicit form of the unitary operators in the decomposition of Eq. 3. While for a general qudit observable $\hat{X}$ the structure of the resulting unitary $\hat{W}_{X}$ may be complicated, the situation simplifies for spin observables. The relevant operator for our case study is the single-spin $\hat{S}_{z}$ operator. In the general spin-$S$ case, the matrix elements with respect to the eigenbasis of the $\hat{S}_{z}$ operator take the simple form
$\big{\langle}m\,\big{|}\,\mathopen{}\hat{W}_{S_{z}}\,\big{|}\,\mathopen{}m^{\prime}\big{\rangle}=\delta_{m,m^{\prime}}\left(\frac{m}{S}+i\sqrt{1-\frac{m^{2}}{S^{2}}}\right)\,,$ (9)
where $m,m^{\prime}\in\{S,\dots,-S\}$. In particular, for $S=1$, we find in matrix representation
$\hat{S}^{z}=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{pmatrix}\>\>\>\>\Rightarrow\>\>\>\>\hat{W}_{S_{z}}=\begin{pmatrix}1&0&0\\ 0&i&0\\ 0&0&-1\end{pmatrix}\,.$ (10)
The unitaries $\hat{W}_{S_{x}}$ and $\hat{W}_{S_{y}}$ can be obtained by applying the basis change that transforms $\hat{S}^{x}$ and $\hat{S}^{y}$ to $\hat{S}^{z}$, respectively. We note that the controlled phase gate represented by $\hat{W}_{z}$ is generally available on state-of-the-art qudit quantum devices, in some cases even as a native operation [61]. Furthermore, for any many-body spin observable $\hat{X}$, we can calculate $W_{\hat{X}}$ explicitly by using the representation of $\hat{X}$ as a sum of spin operator strings, as shown in Appendix C.
III.2.2 Performance benchmark in the presence of shot noise
In an experimental setting, the expectation values in Eq. (11) as well as the outcome of the Hadamard tests are subject to errors caused by finite measurement statistics. We simulate these shot-noise errors by sampling the observables from the ideal distribution given by the total wave function of the system–ancilla register after the execution of the Hadamard-test circuit. To this end, we assign a total budget of shots for measuring the four probabilities on the right-hand side of Eq. 6 as well as the two expectation values forming the disconnected part of the anti-commutator.
Figure 2 compares the exact value of $C_{12}^{\pm}(0,t)$ (blue solid line) to the one obtained from sampling with finite statistics (blue circles). The shot noise is low enough to allow for an accurate measurement of the two-time anti-commutator (commutator) already for a moderate number of $250$ ($2000$) shots for each expectation value, resulting in a total of $1500$ ($8000$) shots per point.
III.3 Comparison to protocols based on linear response
Figure 3: Bias–variance tradeoff in the linear-response measurement of dynamical correlations. With increasing perturbation strength $\lambda$, the linear-response protocol incurs systematic errors in the two-time (a) anti-commutator and (b) commutator due to nonlinear effects. Concomitantly, the statistical error (standard deviation of a single measurement) of the (c) anti-commutator and (d) commutator decreases, requiring a trade-off between accuracy and signal-to-noise ratio. By contrast, the Hadamard-test protocol exhibits consistently lower statistical errors and does not introduce intrinsic systematic sources of error.
In this section, we compare our Hadamard-test protocol with other protocols for measuring dynamical correlation functions. Many such protocols are specific to certain platforms or observables. To keep our analysis general, we focus the following discussion on protocols based on linear response theory [18, 19]. In fact, a large class of well-established protocols for measuring dynamical correlations, including non-invasive (or weak) measurements [44], can be phrased in an implementation-independent way using this framework [25].
III.3.1 (Non-)Hermitian linear-response protocol
As well known in linear response theory, the response of the system to a weak (Hermitian) perturbation gives access to the two-time commutator according to Kubo’s formula [18, 19]. More recently, measuring the two-time anti-commutator by probing the linear response to a non-Hermitian perturbation was proposed [62, 63, 25]. Here, we employ such linear-response protocols to extract the response function directly in the time domain and benchmark the result against our Hadamard-test protocol.
Specifically, to measure the two-time (anti-)commutator $C_{ij}^{\pm}(t_{1},t_{2})$, the system is first evolved under the unperturbed Hamiltonian $\smash{\hat{H}_{0}}$ to the time $t_{1}$. Then, a Hermitian (anti-Hermitian) perturbation of strength $0<\lambda\ll 1$ by the operator $\hat{S}_{j}^{z}$ is applied in the form of a short rectangular pulse of duration $\Delta t$, during which the system evolves under the perturbed Hamiltonian $\hat{H}_{\lambda}=\hat{H}_{0}-\lambda\hbar J_{xy}\hat{S}_{j}^{z}$ (in the non-Hermitian case, we have $\lambda\to i\lambda$). After the perturbation is released, the system is evolved under the unperturbed Hamiltonian $\smash{\hat{H}_{0}}$ up to time $t_{2}$, and the expectation value of the observable $\hat{S}_{i}^{z}$ is measured, yielding the two-time (anti-)commutator via
$C^{\pm}_{ij,\,\mathrm{LR}}(t_{1},t_{2})$ $=\frac{1}{\lambda J_{xy}\Delta t}\big{[}\big{\langle}\hat{S}^{z}_{i}(t_{2})\big{\rangle}_{\pm}-\big{\langle}\hat{S}^{z}_{i}(t_{2})\big{\rangle}\big{]}\,.$ (11)
[…]
III.3.2 Finite statistics and bias–variance trade-off
As depicted in Fig. 2, the linear-response protocol accurately reproduces the time trace of the dynamical correlations for a sufficiently weak perturbation and infinite measurement statistics (red solid line). An advantage of our Hadamard-test protocol over the linear-response method becomes apparent if one takes statistical errors due to shot noise into account. The red circles show the sampled values of the anti-commutator (commutator) using $750$ ($6000$) shots for each expectation value in Eq. 11, corresponding to a total of $1500$ ($12000$) shots per point, which is comparable to the total budget used for the Hadamard-test protocol  [^1] In the non-Hermitian case, we reduce the number of shots according to the decrease of the norm of the state.. One can see that the Hadamard-test protocol yields a lower statistical variance and therefore a higher signal-to-noise ratio.
In order to benchmark the relative performance of the Hadamard-test protocol and the linear-response protocol, we investigate the two figures of merit shown in Fig. 3: the relative error
$\mathcal{R}_{ij}^{\pm}=\frac{\int_{0}^{t}\mathop{}\!\mathrm{d}t^{\prime}\,|C^{\pm}_{ij}(0,t^{\prime})-C^{\pm}_{ij,\,\text{LR}}(0,t^{\prime})|^{2}}{\int_{0}^{t}\mathop{}\!\mathrm{d}t^{\prime}\,|C^{\pm}_{ij}(0,t^{\prime})|^{2}}\,,$ (12)
measuring the systematic error of the linear-response protocol due to non-linear effects, and the time average of the standard deviation,
$\Delta C_{ij}^{\pm}=\frac{1}{t}\int_{0}^{t}\mathop{}\!\mathrm{d}t^{\prime}\Delta C_{ij}^{\pm}(0,t^{\prime})\,,$ (13)
measuring the statistical errors due to shot noise.
The linear-response protocol generally requires a trade-off between accuracy and signal-to-noise ratio. On the one hand, the perturbation should be sufficiently weak that the response is in the linear regime, where nonlinear contributions to the response are negligible. On the other hand, it should be sufficiently strong that the signal is resolvable against the background noise. As can be seen in Fig. 3, the error grows with increasing $\lambda$, while the variance shrinks and vice versa. This behavior can be viewed as a manifestation of the bias–variance tradeo-ff, which is well-known in statistics and machine learning. Importantly, there is no such trade-off in the case of our Hadamard-test protocol. Furthermore, in the investigated quench scenario, the Hadamard-test protocol outperforms the linear-response protocol both in terms of accuracy and signal-to-noise ratio.
III.3.3 Experimental resources
A fair comparison between our Hadamard-test protocol and the linear-response protocol also requires a discussion of the experimental resources required for their implementation. In what follows, we discuss the cases of measuring the unequal-time commutator and anti-commutator separately.
For the commutator, the linear-response protocol requires a standard Hermitian perturbation, which either may be single-body, as in the case of spin–spin dynamical correlations, or may involve few-body interactions. Importantly, the protocol typically admits ancilla-free realizations, where the perturbation is exerted directly on the system, e.g., in the form of a rapid change in the couplings in the Hamiltonian or as a unitary gate in a quantum circuit. By contrast, the protocol based on Hadamard tests requires an ancillary qubit and two entangling operations between this qubit and the system. Thus, depending on the device and scenario at hand, a decision has to be made between additional errors from shot noise and bias (for the linear-response protocol; see Fig. 3) and potential errors from faulty gates (for the Hadamard-test protocol). One way of circumventing the need for a physical ancilla for the Hadamard test is to use qudits with the number of controllable levels equal to double the local Hilbert space dimension of the system we want to simulate (e.g., $d=6$ in the example above).
In our benchmarks of the anti-commutator, we have considered an implementation-agnostic “ideal” non-Hermitian linear-response measurement, where a non-Hermitian perturbation directly acts on the system degrees of freedom. However, physical realizations of this scenario typically incur additional experimental overhead, e.g., entangling the system with an ancilla and performing postselected measurements on the ancilla state [25]. Therefore, engineering an (approximate) effective non-Hermitian Hamiltonian requires resources similar to those of the Hadamard-test protocol.
Finally, we turn our attention to potential near-term experimental realizations of our Hadamard-test protocol for dynamical correlations of general observables. For systems with a large Hilbert space, implementing a general multi-qudit observable as a sum of unitaries presents significant practical challenges. While the decomposition in Eq. 3 applies to any observable, the resulting unitary may have a complicated structure (unlike for spin observables; see Section III.2.1). In this case, the gate complexity may be reduced by resorting to approximate decompositions, e.g., based on the LCU procedure [54], where one can trade off lower circuit depth for a more accurate approximation. While in typical applications the circuit depth is dominated by the number of gates required to implement the system’s time evolution, the added gate complexity due to the (approximate) implementation of the controlled unitaries may be of concern on noisy intermediate-scale quantum (NISQ) devices: In the presence of circuit noise, the maximum circuit depth is restricted by the finite coherence time. This may require lowering the approximation accuracy of the controlled unitaries, introducing systematic errors, and it may impose limits on the times $t_{1}$ and $t_{2}$ of the dynamical correlations that can reliably be measured using the Hadamard test protocol. In situations where this effect is identified as a limiting factor, the linear-response protocol may provide a viable alternative since the implementation of the perturbation is typically favorable in terms of gate complexity.
IV Conclusion
In this work, we proposed an algorithm for measuring dynamical correlation functions on qudit quantum hardware. Key to the scheme is the decomposition of general (multi-)qudit observables into unitaries, which can be exploited to represent dynamical correlations as the sum of quantities accessible via Hadamard tests. We numerically benchmarked the performance of this protocol with finite measurement statistics using the example of a spin-1 XXZ model after a quench. Notably, in this scenario the Hadamard-test protocol outperforms established protocols based on linear response theory in terms of both accuracy and the number of shots needed to distinguish the signal from noise.
The Hadamard-test protocol for dynamical correlation functions can be extended in a straightforward way to measuring higher-order temporal correlations of the form $\langle A_{1}(t_{1})\cdots A_{n}(t_{n})\rangle$. Such a generalization of our protocol would require only the implementation of deeper circuits and a higher number of controlled operations, while the structure of the circuit and the usage of the decomposition of Hermitian operators into a sum of unitaries remain the same.
Finally, we emphasize that our Hadamard-test protocol for measuring dynamical correlations in qudit quantum systems can be used as a subroutine in existing algorithms with wide-ranging applications, ranging from thermalization in many-body systems [24, 25, 65, 53] to the implementation of variational optimization protocols for the simulation of equilibrium and non-equilibrium properties of many-body systems [40, 36, 41]. Moreover, the protocol is designed such that it can be implemented on presently available quantum simulation platforms, including trapped ions, cold atoms, and superconducting qubits.
V Acknowledgements
The ICFO group acknowledges support from: ERC AdG NOQIA; MCIN/AEI [PGC2018-0910.13039/501100011033,  CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, the Plan National STAMEENA PID2022-139099NB-I00 project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI]; QUANTERA MAQS PCI2019-111828-2);  QUANTERA DYNAMITE PCI2022-132919 (QuantERA II Programme co-funded by the European Union’s Horizon 2020 program under Grant Agreement No. 101017733), Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and the European Union through the Recovery, Transformation, and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-1-0013); EU Quantum Flagship (PASQuanS2.1, 101113690); EU Horizon 2020 FET-OPEN OPTOlogic (Grant No. 899794); the EU Horizon Europe Program (Grant Agreement No. 101080086 — NeQST), ICFO Internal “QuantumGaudi” project; the European Union’s Horizon 2020 program under Marie Sklodowska-Curie Grant Agreement No. 847648; and “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID No. 100010434): CF/BQ/PR23/11980043.
This project received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101080086 NeQST, the European Union under NextGenerationEU PRIN 2022 Prot. n. 2022ATM8FY (CUP: E53D23002240006), the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant No. R20PEX7Y3A), the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.4 - Call for tender No. 1031 of the Italian Ministry for University and Research funded by the European Union – NextGenerationEU (Project No. CN_00000013), and Project DYNAMITE QUANTERA2_00056 funded by the Ministry of University and Research through the ERANET COFUND QuantERA II – 2021 call and co-funded by the European Union (H2020, Grant Agreement No. 101017733). This work benefited from Q@TN, the joint laboratory between the University of Trento, Fondazione Bruno Kessler (FBK), National Institute for Nuclear Physics (INFN), and National Research Council (CNR). We acknowledge support from Provincia Autonoma di Trento. P.P.P. also acknowledges support from the “Secretaria d’Universitats i Recerca del Departament de Recerca i Universitats de la Generalitat de Catalunya” under Grant No. 2024 FI-3 00390, as well as the European Social Fund Plus.
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or the Italian Ministry of University and Research. Neither the European Union nor the granting authority can be held responsible for them.
Data Availability
The data set used in this paper is available on Zenodo [66].
Appendix A Outcome of the quantum circuit implementing the Hadamard-test protocol
In this appendix, we derive Eq. 6 by calculating the state of the system after applying each quantum gate in Fig. 1. The system–ancilla register is initialized in the state
$|\Psi^{\alpha}_{0}\rangle=\frac{1}{\sqrt{2}}(|0\rangle+e^{i\alpha}|1\rangle)\otimes|\psi_{0}\rangle$ (14)
with $\alpha\in\{0,\frac{\pi}{2}\}$, depending on whether we want to measure the unequal-time anti-commutator or commutator. After flipping the ancilla, evolving the state up to time $t_{1}$, and applying the first controlled operation $\hat{V}_{A}$, we obtain the intermediate state
$\frac{1}{\sqrt{2}}e^{i\alpha}|0\rangle\otimes\hat{U}(t_{1})|\psi_{0}\rangle+\frac{1}{\sqrt{2}}|1\rangle\otimes\hat{V}_{A}\hat{U}(t_{1})|\psi_{0}\rangle\,.$ (15)
Next, we apply the second ancilla flip, evolve the register up to time $t_{2}$, and apply the second controlled operation $\hat{V}_{B}$, yielding
$\begin{split}\frac{1}{\sqrt{2}}|0\rangle&\otimes\hat{U}(t_{2}-t_{1})\hat{V}_{A}\hat{U}(t_{1})|\psi_{0}\rangle\\ +\frac{1}{\sqrt{2}}e^{i\alpha}|1\rangle&\otimes\hat{V}_{B}\hat{U}(t_{2})|\psi_{0}\rangle\,.\end{split}$ (16)
Finally, after performing the basis transformation on the ancilla via the Hadamard gate, the final state before the measurement reads
$\begin{split}&|\Psi_{V_{A}V_{B}}^{\alpha}(t_{1},t_{2})\rangle\\ =&|0\rangle\otimes\left[\frac{1}{2}\hat{U}(t_{2}-t_{1})\hat{V}_{A}\hat{U}(t_{1})|\psi_{0}\rangle+\frac{1}{2}e^{i\alpha}\hat{V}_{B}\hat{U}(t_{2})|\psi_{0}\rangle\right]\\ +&|1\rangle\otimes\left[\frac{1}{2}\hat{U}(t_{2}-t_{1})\hat{V}_{A}U(t_{1})|\psi_{0}\rangle-\frac{1}{2}e^{i\alpha}\hat{V}_{B}\hat{U}(t_{2})|\psi_{0}\rangle\right]\,.\end{split}$ (17)
The probability for the ancillary qubit to be in the computational state $|0\rangle$ is therefore given by
$2\left|\left\langle 0\middle|\mathopen{}\Psi_{V_{A}V_{B}}^{\alpha}(t_{1},t_{2})\right\rangle\right|^{2}-1\\ =\frac{1}{2}e^{-i\alpha}\big{\langle}\psi_{0}\,\big{|}\,\mathopen{}\hat{U}^{\dagger}(t_{2})\hat{V}^{\dagger}_{B}\hat{U}(t_{2}-t_{1})\hat{V}_{A}\hat{U}(t_{1})\,\big{|}\,\mathopen{}\psi_{0}\big{\rangle}+\text{c.c.}\,,$ (18)
where c.c. denotes the complex conjugate. By setting $\alpha=\alpha_{+}\coloneqq{}0$ and $\alpha=\alpha_{-}\coloneqq{}\pi/2$, we get the real and the imaginary parts of the dynamical correlation function of $\hat{V}_{A}$ and $\hat{V}^{\dagger}_{B}$, respectively:
$1\displaystyle 2P_{V_{A}V_{B}}^{+}(t_{1},t_{2})-$ 1 $=\frac{1}{2}\big{\langle}\hat{V}_{A}^{\dagger}(t_{1})\hat{V}_{B}(t_{2})+\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{2})\big{\rangle}$
$=\big{\langle}\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{1})\big{\rangle}\,,$ (19a)
$1\displaystyle 2P_{V_{A}V_{B}}^{-}(t_{1},t_{2})-$ 1 $=\frac{1}{2}i\big{\langle}\hat{V}_{A}^{\dagger}(t_{1})\hat{V}_{B}(t_{2})-\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{1})\big{\rangle}$
$=\operatorname{Im}\big{\langle}\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{1})\big{\rangle}\,.$ (19b)
By means of the decomposition in Eq. 5, we can express the two-time (anti-)commutator as
$C_{AB}^{\pm}$ $=\frac{1}{4}\lVert\hat{A}\rVert\lVert\hat{B}\rVert\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\big{\langle}e^{i\alpha_{\pm}}\hat{V}_{A}\hat{V}_{B}^{\prime}+e^{-i\alpha_{\pm}}\hat{V}_{B}^{\prime}\hat{V}_{A}\big{\rangle}$
$=\frac{1}{4}\lVert\hat{A}\rVert\lVert\hat{B}\rVert\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\big{\langle}e^{-i\alpha_{\pm}}\hat{V}_{B}^{\prime\,\dagger}\hat{V}_{A}+e^{i\alpha_{\pm}}\hat{V}_{A}^{\dagger}\hat{V}_{B}^{\prime}\big{\rangle}$
$=\frac{1}{2}\lVert\hat{A}\rVert\lVert\hat{B}\rVert\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\big{(}\langle\hat{V}_{B}^{\prime\,\dagger}\hat{V}_{A}\rangle\cos\alpha_{\pm}$
[…]
where in the second line we have rearranged the terms in the sum. Combining this expression with Eq. 19 finally yields the desired result in Eq. 6.
Appendix B Variance of the Hadamard test
We proceed with the proof of Eq. 7. To this end, we rewrite Eq. 6 by expressing the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ of measuring $|0\rangle$ on the ancilla as the expectation value of the projector $\hat{\Pi}_{0}=|0\rangle\langle 0|\otimes\mathds{1}$:
$C_{AB}^{\pm}(t_{1},t_{2})=\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\\ \times\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\Big{(}\big{\langle}\Psi_{V_{A}V_{B}^{\prime}}^{\pm}(t_{1},t_{2})\,\big{|}\,\mathopen{}\hat{\Pi}_{0}\,\big{|}\,\mathopen{}\Psi_{V_{A}V_{B}^{\prime}}^{\pm}(t_{1},t_{2})\big{\rangle}-\frac{1}{2}\Big{)}\,.$ (21)
Since the measurements of the four Hadamard-test configurations with controlled unitaries $\hat{V}_{A}\in\{\hat{W}_{A},\hat{W}^{\dagger}_{A}\}$ and $\hat{V}_{B}\in\{\hat{W}_{B},\hat{W}^{\dagger}_{B}\}$ are independent, standard propagation of variance yields [67]
$\operatorname{Var}[C^{\pm}_{AB}(t_{1},t_{2})]=\big{\lVert}\hat{A}\big{\rVert}^{2}\big{\lVert}\hat{B}\big{\rVert}^{2}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\operatorname{Var}[\hat{\Pi}_{0}]^{\pm}_{V_{A}V_{B}^{\prime}}(t_{1},t_{2})\,.$ (22)
If the Hadamard test with controlled unitaries $\hat{V}_{A}$ and $\hat{V}_{B}$ is repeated $M_{V_{A}V_{B}}$ times, the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ can be estimated with variance
$\operatorname{Var}[\hat{\Pi}_{0}]^{\pm}_{V_{A}V_{B}}(t_{1},t_{2})\\ \begin{split}=\frac{1}{M_{V_{A}V_{B}}}\Big{[}&\big{\langle}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\,\big{|}\,\mathopen{}\hat{\Pi}_{0}^{2}\,\big{|}\,\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\big{\rangle}\\ &-\big{\langle}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\,\big{|}\,\mathopen{}\hat{\Pi}_{0}\,\big{|}\,\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\big{\rangle}^{2}\Big{]}\,.\end{split}$ (23)
For the projector $\hat{\Pi}_{0}$, the simple identity $\hat{\Pi}_{0}^{2}=\hat{\Pi}_{0}$ holds, and thus for an arbitrary state
$\langle\hat{\Pi}_{0}^{2}\rangle-\langle\hat{\Pi}_{0}\rangle^{2}=\langle\hat{\Pi}_{0}\rangle\big{(}1-\langle\hat{\Pi}_{0}\rangle\big{)}\,.$ (24)
Finally, identifying the expectation value $\langle\hat{\Pi}_{0}\rangle$ with the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})$ concludes the proof.
In order to arrive from Eq. 7 at the a priori upper bound $\operatorname{Var}[C^{\pm}_{AB}(t_{1},t_{2})]\leq M^{-1}\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$, we used the fact that the function $f(x)=x(1-x)$ has a maximum value $f(x=1/2)=1/4$ in the interval $[0,1]$.
Appendix C Decomposition for many-body observables
Typical $n$-spin observables are given by strings of the form
$\hat{X}=\hat{S}^{\alpha_{1}}_{1}\otimes\cdots\otimes\hat{S}^{\alpha_{n}}_{n}\,,$ (25)
where each $\hat{S}^{\alpha_{k}}_{k}$, with $\alpha_{k}\in\{x,y,z\}$, is a spin operator. The unitary $\hat{W}_{X}$ can then be calculated using the unitary decomposition of each of the (one-body) spin operators:
$\hat{S}^{\alpha_{k}}_{k}=\frac{1}{2}\lVert\hat{S}^{\alpha_{k}}_{k}\rVert(\hat{W}_{\alpha_{k}}+\hat{W}^{\dagger}_{\alpha_{k}})\,.$ (26)
Inserting Eq. (26) into Eq. (25) and using the definition of $\hat{W}_{X}$, we get
$\hat{W}_{X}=\hskip 10.00002pt\sum_{\mathclap{\left\{V_{k}=W_{\alpha_{k}},W_{\alpha_{k}}^{\dagger}\right\}}}\hat{V}_{1}\otimes\cdots\otimes\hat{V}_{n}\,.$ (27)
In this case, the experimental implementation of the controlled unitaries in the Hadamard-test simplifies since we have to deal with only a product of two-body operations. However, the number of Hadamard tests to be performed increases exponentially with the length of the string in Eq. (25).
For a general multi-qudit observable, representations similar to strings of one-qudit observables can be employed to simplify the experimental implementation of controlled operations in the Hadamard test. In practice, however, an approximate decomposition such as the one resulting from the LCU procedure [54] may be more efficient in terms of circuit depth for implementations on NISQ devices.